You are tasked with investigating whether fake news media platform and personality are independent. Consider the table below to help you with your task.
Celebrity. C/ Politician .P/ Sport Star .SS/ Total
Social Media .SM/ 1800 700 515 3015
Traditional Media. TM/ 485 350 150 985
Total 2285 1050 775 4000

Consider the statements below. Assume the level of significance is 1%
A. We reject the null hypothesis
B. We do not reject the null hypothesis
C. Media platform and personality are independent

Which statements are correct?
1. Only A
2. Only B
3. Only C
4. A and C Only
5. B and C only
1
Expert's answer
2020-10-01T10:14:23-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 & C & P & SS & Total \\ \hline
 SM & 1800 & 700 & 515 & 3015 \\
 \hdashline
 TM & 485 & 350 & 150 & 985 \\
 \hdashline
 Total & 2285 & 1050 & 665 & 4000
\end{array}"

The expected values are computed in terms of row and column totals. In fact, the formula is



"E_{ij}=\dfrac{R_i\times C_j}{T},"

where "R_i" corresponds to the total sum of elements in row "i," "C_j" corresponds to the total sum of elements in column "j," and "T" is the grand total.



"E_{11}=\dfrac{3015\times 2285}{4000}=1722.31875""E_{12}=\dfrac{3015\times 1050}{4000}=791.4375"




"E_{13}=\dfrac{3015\times 665}{4000}=501.24375""E_{21}=\dfrac{985\times 2285}{4000}=562.68125""E_{22}=\dfrac{985\times 1050}{4000}=258.5625""E_{23}=\dfrac{985\times 665}{4000}=163.75625"

The table below shows the calculations to obtain the table with expected values:



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 Expected\ Values & C & P & SS & Total \\ \hline
 SM & 1722.32 & 791.44& 501.24 & 3015 \\
 \hdashline
 TM & 562.68 & 258.56 & 163.76 & 985 \\
 \hdashline
 Total & 2285 & 1050 & 665 & 4000
\end{array}"

Based on the observed and expected values, the squared distances can be computed according to the following formula: "\dfrac{(E-O)^2}{E}." The table with squared distances is shown below:



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 Squared\ Distances & C & P & SS \\ \hline
 SM & 3.504 & 10.564 & 0.378 \\
 \hdashline
 TM & 10.724 & 32.336 & 1.156
\end{array}"

The following null and alternative hypotheses need to be tested:

"H_0:" The two variables are independent.

"H_1:" The two variables are dependent.

This corresponds to a Chi-Square test of independence.

Based on the information provided, the significance level is "\alpha=0.05," the number of degrees of freedom is "df=(2-1)\times(3-1)=2," so then the rejection region for this test is "R=\{\chi^2:\chi^2>5.991\}."

The Chi-Squared statistic is computed as follows:



"\chi^2=3.504+10.724+10.564+32.336+""+0.378+1.156=58.661"

Since it is observed that "\chi^2=58.661>5.991=\chi_c^2," it is then concluded that the null hypothesis is rejected.Therefore, there is enough evidence to claim that the two variables are dependent, at the 0.05 significance level.


1. Only A statement is correct.